The base of an isosceles triangle is 8 feet long. If the altitude is 6 feet long and increasing 3 inches per minute, at what rate are the base angles changing?


1
Expert's answer
2021-01-26T02:37:07-0500

Considering the isosceles triangle,

"\text{where h is the altitude}\\
tan(\theta)=\frac{h}{4}\\
\text{we differentiate both sides}\\
sec^2\theta \frac{d\theta}{dt}=\frac{1}{4}\frac{dh}{dt}\\
h=6, \frac{dh}{dt}=3\\
\text{from trigonometric identities}\\
sec^2\theta \frac{d\theta}{dt}=1+tan^2\theta\\
1+(\frac{6}{4})^2=\frac{13}{4}\\
\frac{13}{4} \frac{d\theta}{dt}=\frac{1}{4}\times3\\
13 \frac{d\theta}{dt}=3\\
\frac{d\theta}{dt}=\frac{3}{13}\text{radian per minutes}"